OpenStudy (anonymous):

A family has two cars. The first car has a fuel efficiency of 20 miles per gallon of gas and the second has a fuel efficiency of 25 miles per gallon of gas. During one particular week, the two cars went a combined total of 1025 miles, for a total gas consumption of 45 gallons. How many gallons were consumed by each of the two cars that week?

OpenStudy (dumbcow):

x + y = 45 -------> y = 45-x 20x + 25y = 1025 solve with substitution